OpenStudy (anonymous):

if 4 shirts cost $76, what is the cost of 2 shirts?

OpenStudy (anonymous):

76/2 = 38 So $38

OpenStudy (anonymous):

thx ur the best

OpenStudy (anonymous):

so just divide each shirt number right?

OpenStudy (anonymous):

ttyl

OpenStudy (anonymous):

You have to cut 4 in half, because theyre asking for the price of 2. And so you do 76 divided by 2. Your answer is 38

OpenStudy (texaschic101):

4/76 = 2/x -- (4 shirts for 76 = 2 shirts for x) cross multiply (4)(x) = (76)(2) 4x = 152 x = 152/4 x = 38 yep..thats correct

OpenStudy (anonymous):

I don't think this person is learning how to cross multiply yet. Keep it simple :)

OpenStudy (texaschic101):

your probably right....thx

OpenStudy (anonymous):

No problem